# Phase space and Data analyses of a non-minimally coupled scalar field system with decaying dark energy model

### v1

@@ Abstract
We investigate a class of scalar field dark energy models non-minimally coupled to gravity, characterized by a double exponential potential and parameterized coupling $\xi$. First, we study the cosmological dynamics for a recently proposed descending dark energy model, namely, Q-SC-CDM. Initially, we choose two distinct values of $\xi$: 0.1 and 0.5. In case of $\xi=0.1$, the evolution of the universe is split up into three different phases: decelerated expansion (early time), accelerated expansion (late-time) and slow-contraction (future era), and provide Big Crunch Singularity at distant future. However, in case of $\xi=0.5$, the phase of slow-contraction vanishes, cosmic acceleration is obtained for longer period, and the universe gets de-Sitter expansion at distant future. It is remarkable to see that the Big Crunch Singularity is redundant in the later case. Second, we perform the phase space analysis for the model under consideration. Our investigation brings new asymptotic regimes and finds stable de-Sitter solution. We obtain phase portrait, where all trajectories move towards stable attractor point. Finally, fixing $\xi$ to two representative values: $\xi=0.1$ and $\xi=0.05$, we perform a comprehensive Bayesian analysis using recent late-time cosmological observations, including Cosmic Chronometers (CC), Type Ia Supernovae (Pantheon+ and DES-SN5YR), and Baryon Acoustic Oscillation (DESI DR2) data. Our results demonstrate that both models yield constraints on key cosmological parameters—$\Omega_{0m}$, $H_{0}$ and the sound horizon $r_{d}$—that are consistent with $\Lambda$CDM within 68% confidence level, yet exhibit mild tension with Pantheon+ measurements. We analyze the evolution of the effective equation of state, showing that the model transitions from a stiff matter phase at high redshift to a dark energy dominated phase with effective equation of state $<-0.5$ at late times. Additionally, we employ the $Om(z)$ diagnostic to distinguish our model from $\Lambda$CDM, finding minimal deviation up to redshift $z\sim 2$. Statistical model comparison using Akaike (AIC) and Bayesian (BIC) Information Criteria reveals moderate support for the model with $\xi=0.1$, though $\Lambda$CDM remains statistically preferred. Overall, our findings suggest that non-minimally coupled scalar field models remain viable alternatives to the cosmological constant, providing competitive fits to late-time data while allowing for dynamical dark energy behavior.
@@ I Introduction
In the standard quintessence scenario, the energy density of the scalar field minimally coupled to gravity mimics the effective cosmological constant. The detailed evolution is obviously dependent on a particular form of the potential, but the $\phi^{2}$ contribution can be regarded as a leading order term of expansion of the potential $V(\phi)$. We incorporate the non-minimal coupling constant to the quintessence scenario. Lykkas and Perivolaropoulos demonstrate that certain values of the non-minimal coupling parameter can prevent the cosmic apocalypse singularity in scalar-tensor quintessence with a linear potential [18]. For many years, researchers have examined the scalar tensor theories, which have a scalar field that is non-minimally coupled (NMC) to gravity. In order to reconcile Mach’s principle with general relativity, the first well-known instance of non-minimal coupling à $l$a Brans-Dicke theory was put forth in 1961 [19]. According to this theory, the gravitational constant is replaced by a scalar field $\phi$ entering the action in a particular combination with a Riemannian curvature as $\phi^{2}R$. Other types of scalar-tensor action were studied after the Brans-Dicke proposal: a well-known example of a NMC scalar field system is given by $F(\phi)R$ coupling with $F(\phi)=1-\xi\phi^{2}$. Such a theory has quite rich cosmic dynamics that need consideration. For a recent breakthrough, a NMC Higgs field because of a large coupling $\xi$ could result in a successful inflation, which is otherwise impossible [20]. The construction of a dark energy model is very interested in NMC scalar field systems because of their unique characteristics [21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47]. Non-minimal coupling, for example, may enable phantom crossover and result in cosmic scaling solutions that are relevant to dark energy models. The general characteristics of a NMC system with $F(\phi)=1-\xi\phi^{2}$ are phantom scaling solutions [48]. In the literature, two values of the coupling constant are frequently considered: $\xi=0$ (minimal coupling) and $\xi=1/6$ (conformal coupling). However, in many cases, $|\xi|\gg 1$ (strong coupling) has also been used [49, 50, 51, 52, 53, 54]. Recently, the value of $\xi$ was estimated from a distant supernova [55].
@@
Dynamical system theory has been widely applied in cosmology to provide a broad understanding of dynamics for a variety of cosmological models. This approach has the benefit of having some sort of “machines” that uses a straightforward programmed algorithm to derive asymptotic solutions. In order to rewrite the original system as a system of first-order differential equations, a new set of variables must be introduced. This approach allows for the derivation of asymptotic solutions, and stability can be assessed using a straightforward programmed algorithm. Traditionally, physical cosmological solutions are derived from the combination of phase portraits and the stability of critical points. The ability to display any system trajectory that is permissible for every initial condition is the primary benefit of dynamical system analysis. Consequently, generic pathways to the accelerating phase (the de-Sitter attractor) can be categorized. This paper aims to explore the cosmological dynamics of the Quintessence-Driven Slow-Contraction Cold Dark Matter (Q-SC-CDM) model and to examine its stationary points and stability. The Q-SC-CDM model was initially introduced by C. Andrei et al. in the framework of decaying dark energy and the end of cosmic expansion [56, 57]. In this research, we examine the cosmological dynamics of a NMC scalar field model with double exponential potential and a specific form of $F(\phi)$. We limit ourselves to the functional form $F(\phi)=1-\xi\phi^{2}$. We shall analyze the cosmological dynamics of the model under consideration and discuss the regions of deceleration, cosmic acceleration and slow contraction. Furthermore, we shall conduct a phase space analysis of the Q-SC-CDM model, and investigate the existence of a stable late-time de-Sitter attractor solution that corresponds to an EOS of $-1$ and an energy density parameter $\Omega_{\phi}=1$. The paper is organized as follows. In section II, we discuss the equations of motion for a NMC scalar field system in a homogeneous and isotropic flat Friedmann–Lemaître–Robertson–Walker (FLRW) universe. The subsection II.1 is devoted to the cosmological dynamics of NMC system for different values of $\xi$, and explored the phases of deceleration, cosmic acceleration and contraction of the universe. In section III and subsection III.1, we construct the autonomous system that is useful for the dynamical system analysis, and discuss the stationary points, stability and cosmological importance for Q-SC-CDM model. In section IV, we reformulate the equations of motion to enhance the numerical stability. The sections V and VI are devoted to the data analysis and their results. The conclusion is presented in section VII.
@@ II.1 Cosmological Dynamics of Descending Dark Energy
The decaying dark energy model suggests that dark energy is not a constant but gradually loses energy over time. In such type of model, the energy density of dark energy decreases as the universe evolves, either through a direct decay into other forms of energy or as a function of the scale factor of the universe. This is different from the standard cosmological model, which assumes that dark energy remains constant in its density (described by the cosmological constant $\Lambda$). Recently, a new model of decaying dark energy, namely, Q-SC-CDM was proposed by C. Andrei $et$ $al.$ [56]. The Q-SC-CDM scenario proposes that the universe underwent a period of slow-contraction rather than acceleration. In this case, quintessence acts as the driving force for the contraction and evolves in such a way that it causes the cosmic scale factor to shrink over time. We use the following potential proposed by C. Andrei $et$ $al.$ [56, 57].
@@
Figure 3: The figure is similar to Fig. 2 except $\xi=0.5$. In this case, we get $q=-0.53$, $w_{eff}=-0.72$, $w_{\phi}=-0.98$ and $\Omega_{\phi}=0.70$ at current epoch. For large value of $\xi$ (e.g. $\xi=0.5$), the phase of slow-contraction vanishes, and we obtain cosmic acceleration at current epoch and de-Sitter expansion in distant future.
@@
Let us make some comments on the value of non-minimal coupling $\xi$. As $\xi\rightarrow 0$, the standard quintessence scenario is recovered. In this case, the evolution of the universe is divided into three different phases such as deceleration (past era), acceleration (current epoch) and slow-contraction (future era) [56]. In quintessence models where potentials become negative may give rise to collapse of the universe in future, which is known as Big Crunch Singularity. Such singularity can be avoided in scalar-tensor theories with non-minimal coupling [18]. We numerically check the range of $\xi$ to alleviate the Big Crunch Singularity. For $0\leq\xi\leq 0.3$, this kind of singularity can not be avoided whereas it vanishes for $\xi>0.3$. In what follows let us consider two distinct cases one by one. As we move a little bit higher value of $\xi$ (e.g. $\xi=0.1$, see Fig. 2), we still obtain three different phases as mentioned earlier. However, the regime of slow-contraction is shifted in far future in comparison to the case of $\xi=0$. In other words, the Big Crunch Singularity is delayed for $\xi=0.1$. In case of $\xi=0.5$, the slow-contraction regime does not exist any more, an accelerated expansion is obtained for longer period, and finally the universe gets de-Sitter expansion in distant future, see Fig. 3. Therefore, the Big Crunch Singularity is redundant for large values of non-minimal coupling (e.g. $\xi=0.5$) in NMC scalar field model. The numerical evolution of scale factor $a(t)$, Hubble parameter $H(t)$ and deceleration parameter $q(t)$ of NMC scalar field model are well approximated by $\Lambda$CDM as shown in Fig. 3, where we get an accelerated expansion at current epoch and de-Sitter expansion in far future.
@@ III Phase Space Analysis: Stationary points and their stability
The equations of motion (2), (3), (4) and (5) are reformulated in new variables as a system of differential equations of first order.
@@ III.1 Cosmological Importance
Points $A_{3}$ and $A_{4}$ (Stable Attractor Solution): The fixed points $A_{3}$ and $A_{4}$ are stable attractors, which behave as a stable node with eigenvalues given by equations (III) and (III), respectively. It further shows that the solutions exhibit damped oscillatory behavior along the real parts of the eigenvalues with a stable attraction to the fixed point. At these points, the universe shows accelerating behavior with $\Omega_{\phi}=1$, meaning the scalar field dominates the energy density, while $\Omega_{m}=0$ indicates no matter content. The EoS is $w_{eff}=w_{\phi}=-1$, corresponding to a cosmological constant-like behavior, which is consistent with an accelerated expansion of the universe. The phase portraits provide crucial insight into the attractor behavior for both the points. As shown in Fig. 4, all trajectories approach the attractor point, which validates the stability of the fixed point and its role in describing the late-time behavior of the universe, particularly in the context of dark energy dominance. The stable node is typical of models where dark energy dominates at late times. It corresponds to a cosmological constant or quintessence field behavior, where the universe enters a phase of accelerated expansion. The eigenvalue structure and stability align well with common treatments of scalar field models with exponential potentials.
@@ IV Equations for Data analysis
In this section, we reconstruct the first-order dynamical equations that facilitate the evaluation of the Hubble parameter in our model. In the previous section, we analyzed the qualitative behavior of the system using the dynamical systems framework. However, the equations expressed in that form exhibit numerical instability. Such instability is common in highly non-linear models; therefore, reformulating the system in a slightly different manner can significantly enhance numerical stability. It is important to note that the autonomous system presented earlier can still be used to derive constraints on the model parameters. In the previous works, the parameters for the $k-$essence and quintessence scalar field model have been constrained using the dynamical system stability analysis framework [65, 66, 67]. The reformulation presented here does not alter the physical interpretation or the parameter constraints of the system. Therefore, the redefined variables yield equivalent information. The redefined dimensionless dynamical variables are
@@ V Data Analysis
  • BAO data: The Baryonic Acoustic Oscillation (BAO) dataset, referred to as DESBAO, consists of 7 data points from the Dark Energy Spectroscopic Instrument (DESI) Release II [74], which is an updated and enhanced version of DR1 [73, 76, 77]. We derive the constraints using the full covariance matrix provided in [74]. The relevant observables include $(D_{M},D_{H},D_{V},r_{d})$, corresponding to the comoving angular distance, comoving Hubble distance, spherically averaged distance, and the sound horizon at the drag epoch, respectively. For mathematical definitions of these parameters, we refer readers to Refs. [75, 73].
    In this study, we treat $r_{d}$ as a free parameter, which we constrain using various combinations of datasets. Specifically, we consider two distinct combinations:
@@
    where $k$ is the number of model parameters, $N$ is the total number of data points, and $\mathcal{L}_{\rm max}$ is the maximum likelihood. To assess model preference, we compute the difference in information criteria between the current model and $\Lambda$CDM:
@@ VI Results
Additionally we present the correlation matrix in Fig. 6, generated using Seaborn python package, utilizing the sample chains obtain from MCMC [86]. The matrix is constructed by combining samples from both datasets, as the parameter correlations are found to be qualitatively similar across them. In this matrix, higher positive (or negative) values indicate stronger correlations between the parameters. A positive correlation implies that as one parameter increases, the other tends to increase proportionally, whereas a negative correlation indicates inverse proportionality. For example, $H_{0}$ and $r_{d}$ exhibit a strong negative correlation, implying that an increase in $H_{0}$ corresponds to a decrease in $r_{d}$, and vice versa. A similar inverse correlation is observed between $\Omega_{0m}$ and $r_{d}$. In contrast, the correlation between $\Omega_{0m}$ and $H_{0}$ is weak and also negative. Among the model parameters, $\alpha$ and $\beta$, which define the structure of the potential, are strongly anti-correlated. Furthermore, $\alpha$ is positively correlated with $z_{0}$, the initial value of the field variable $z=\kappa\phi$, while showing negligible correlation with $\beta$. The parameter $A_{0}$, on the other hand, shows negative correlation with both $\alpha$ and $\beta$, suggesting a suppressive effect on the potential parameters.
@@
The qualitative behavior of the models is illustrated in Fig. 7 as a function of redshift $z$, where we plot the fractional energy densities of the scalar field and matter fluid, the EoS of the field, and the effective EoS of the system, using the best-fit parameter values obtained from the MCMC analysis. The evolution of the field energy density indicates that it begins to dominate near the current epoch, with the corresponding EoS of the field approaching $\omega_{\phi}\approx-1$, while the effective EoS satisfies $\omega_{eff}<-0.5$, mimicking the behavior of the $\Lambda$CDM model.
At present, the matter density is approximately $\Omega_{m}\sim 0.3$, and it increases at higher redshifts, reflecting the expected matter-dominated era where the effective EoS tends toward zero. As the system is extrapolated into the asymptotic past, the effective EoS rises toward $\omega_{eff}\to 1$, signaling a stiff matter-like behavior, with $\Omega_{eff}$ correspondingly dominating. It is important to note that, since radiation has not been included in this system, the onset of stiff matter behavior occurs at intermediate redshifts for specific parameter configurations. Inclusion of a radiation component would shift this stiff phase further back into the early universe. However, given that the observational data used in this study are largely limited to redshifts $z\lesssim 3$, the radiation density does not significantly affect our late-time cosmological analysis.
@@
We assess the statistical performance of the proposed models using information criteria and compare them with the standard $\Lambda$CDM model as shown in Fig. 9. In addition, we compute the reduced chi-square statistic, defined as $\chi_{\rm rd}^{2}\equiv\chi_{\rm min}^{2}/\nu$, where $\nu$ represents the number of observational data points minus the number of fitted parameters. This metric provides an important measure of goodness-of-fit. A value of $\chi_{\rm rd}^{2}\ll 1$ typically indicates model overfitting, whereas $\chi_{\rm rd}^{2}\gg 1$ suggests a poor fit. A value near unity implies that the model fits the data well. Our results show that the proposed models yield $\chi_{\rm rd}^{2}\sim 1$, consistent with that of the $\Lambda$CDM model, indicating excellent agreement with the observational data.
Furthermore, we find that the current models yield a $\Delta\rm AIC$ of approximately 5, suggesting moderate support for the models relative to $\Lambda$CDM, though not sufficient for a strong preference. Conversely, the $\Delta\rm BIC>10$ indicates weak support for the current models when penalizing model complexity more strongly. Among the two considered cases, the model with $\xi=0.1$ exhibits a slightly better performance in both AIC and BIC metrics compared to the model with $\xi=0.05$, thus indicating a mild statistical preference.

### v2

# Phase space and data analyses of a nonminimally coupled scalar field system with a decaying dark energy model

@@ Abstract
We investigate a class of scalar field dark energy models non-minimally coupled to gravity, characterized by a double exponential potential and parameterized coupling $\xi$. We study the cosmological dynamics for a recently proposed descending dark energy model, namely, Q-SC-CDM. Initially, we choose distinct values of coupling parameter. For some values of $\xi$, the evolution of the universe is split up into three different phases: decelerated expansion (early time), accelerated expansion (late-time) and slow-contraction (future era), and provide Big Crunch Singularity at distant future. In other scenario, the phase of slow-contraction vanishes, cosmic acceleration is obtained at current epoch, and the universe gets de-Sitter expansion at distant future. It is remarkable to see that the Big Crunch Singularity is redundant in the later case. Next, we investigate the phase space analysis for the model under consideration. Our investigation brings new asymptotic regimes and finds stable de-Sitter solution. Eventually, we perform a comprehensive Bayesian analysis using recent cosmological observations, including Cosmic Chronometers, Type Ia Supernovae (Pantheon+ and DES-SN5YR), and Baryon Acoustic Oscillation (DESI DR2) data. The results demonstrate that the present model yield constraints on key cosmological parameters—$\Omega_{0m}$, $H_{0}$ and the sound horizon $r_{d}$—that are consistent with $\Lambda$CDM within 68% confidence level, yet exhibit mild tension with Pantheon+ measurements. Additionally, we employ the $Om(z)$ diagnostic test, Akaike and Bayesian Information Criteria to distinguish our model from $\Lambda$CDM. The Statistical comparison reveals moderate support for the current model.
@@ I Introduction
In the standard quintessence scenario, the energy density of the scalar field minimally coupled to gravity mimics the effective cosmological constant. The detailed evolution is obviously dependent on a particular form of the potential, but the $\phi^{2}$ contribution can be regarded as a leading order term of expansion of the potential $V(\phi)$. We incorporate the non-minimal coupling constant to the quintessence scenario. Lykkas and Perivolaropoulos demonstrate that certain values of the non-minimal coupling parameter can prevent the cosmic apocalypse singularity in scalar-tensor quintessence with a linear potential [18]. For many years, researchers have examined the scalar tensor theories, which have a scalar field that is non-minimally coupled (NMC) to gravity. In order to reconcile Mach’s principle with general relativity, the first well-known instance of non-minimal coupling à $l$a Brans-Dicke theory was put forth in 1961 [19]. According to this theory, the gravitational constant is replaced by a scalar field $\phi$ entering the action in a particular combination with a Riemannian curvature as $\phi^{2}R$. Other types of scalar-tensor action were studied after the Brans-Dicke proposal: a well-known example of a NMC scalar field system is given by $F(\phi)R$ coupling with $F(\phi)=1-\xi\phi^{2}$. Such a theory has quite rich cosmic dynamics that need consideration. For a recent breakthrough, a NMC Higgs field because of a large coupling $\xi$ could result in a successful inflation, which is otherwise impossible [20]. The construction of a dark energy model is very interested in NMC scalar field systems because of their unique characteristics [21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51]. Non-minimal coupling, for example, may enable phantom crossover and result in cosmic scaling solutions that are relevant to dark energy models. The general characteristics of a NMC system with $F(\phi)=1-\xi\phi^{2}$ are phantom scaling solutions [52]. In the literature, two values of the coupling constant are frequently considered: $\xi=0$ (minimal coupling) and $\xi=1/6$ (conformal coupling). However, in many cases, $|\xi|\gg 1$ (strong coupling) has also been used [53, 54, 55, 56, 57, 58]. Recently, the value of $\xi$ was estimated from a distant supernova [59].
@@
Dynamical system theory has been widely applied in cosmology to provide a broad understanding of dynamics for a variety of cosmological models. This approach has the benefit of having some sort of “machines” that uses a straightforward programmed algorithm to derive asymptotic solutions. In order to rewrite the original system as a system of first-order differential equations, a new set of variables must be introduced. This approach allows for the derivation of asymptotic solutions, and stability can be assessed using a straightforward programmed algorithm. Traditionally, physical cosmological solutions are derived from the combination of phase portraits and the stability of critical points. The ability to display any system trajectory that is permissible for every initial condition is the primary benefit of dynamical system analysis. Consequently, generic pathways to the accelerating phase (de-Sitter attractor) can be categorized. This paper aims to explore the cosmological dynamics of the Quintessence-Driven Slow-Contraction Cold Dark Matter (Q-SC-CDM) model and to examine its stationary points and stability. The Q-SC-CDM model was initially introduced by C. Andrei et al. in the framework of decaying dark energy and the end of cosmic expansion [60, 61]. In this research, we examine the cosmological dynamics of a NMC scalar field model with double exponential potential and a specific form of $F(\phi)$. We limit ourselves to the functional form $F(\phi)=1-\xi\phi^{2}$. We shall analyze the cosmological dynamics of the model under consideration and discuss the regions of deceleration, cosmic acceleration and slow-contraction. Furthermore, we shall conduct a phase space analysis of the Q-SC-CDM model, and investigate the existence of a stable late-time de-Sitter attractor solution that corresponds to an EoS of $-1$ and an energy density parameter $\Omega_{\phi}=1$. The paper is organized as follows. In section II, we discuss the equations of motion for a NMC scalar field system in a homogeneous and isotropic flat Friedmann–Lemaître–Robertson–Walker (FLRW) universe. The subsection II.1 is devoted to the cosmological dynamics of NMC system for different values of $\xi$, and explored the phases of deceleration, cosmic acceleration and contraction of the universe. In section III and subsection III.1, we construct the autonomous system that is useful for the dynamical system analysis, and discuss the stationary points, stability and cosmological importance for Q-SC-CDM model. In section IV, we reformulate the equations of motion to enhance the numerical stability. The sections V and VI are devoted to the data analysis and their results and discussion, respectively. The conclusion is presented in section VII.
@@ II.1 Cosmological Dynamics of Descending Dark Energy
The decaying dark energy model suggests that dark energy is not a constant but gradually loses energy over time. In such a model, the energy density of dark energy decreases as the universe evolves, either through a direct decay into other forms of energy or as a function of the scale factor of the universe. This is different from the standard cosmological model, which assumes that dark energy remains constant in its density (described by the cosmological constant $\Lambda$). Recently, a new model of decaying dark energy, namely, Q-SC-CDM was proposed by C. Andrei $et$ $al.$ [60]. The Q-SC-CDM scenario proposes that the universe underwent a period of slow-contraction rather than acceleration. In this case, quintessence acts as the driving force for the contraction and evolves in such a way that it causes the cosmic scale factor to shrink over time. We use the following potential proposed by C. Andrei $et$ $al.$ [60, 61].
@@
Figure 3: The figure is similar to Fig. 2 except $\xi=0.5$. In this case, we get $q=-0.53$, $w_{eff}=-0.72$, $w_{\phi}=-0.98$ and $\Omega_{\phi}=0.70$ at current epoch. Moreover, the phase of slow-contraction vanishes, and we obtain cosmic acceleration at current epoch and de-Sitter expansion in distant future.
@@
Let us make some comments on the value of non-minimal coupling $\xi$. As $\xi\rightarrow 0$, the standard quintessence scenario is recovered. In this case, the evolution of the universe is divided into three different phases such as deceleration (past era), acceleration (current epoch) and slow-contraction (future era) [60]. In quintessence models where potentials become negative may give rise to collapse of the universe in future, which is known as Big Crunch Singularity. Such singularity can be avoided in scalar-tensor theories with non-minimal coupling [18]. We numerically check the range of $\xi$ to alleviate the Big Crunch Singularity. For $0.20\leq\xi<$ 0.2, this kind of singularity can not be avoided whereas it vanishes for $\xi\geq 0.2$. In what follows let us consider two distinct cases one by one. As we move a little bit higher value of $\xi$ (e.g. $\xi=0.1$, see Fig. 2), we still obtain three different phases as mentioned earlier. However, the regime of slow-contraction is shifted in far future in comparison to the case of $\xi=0$ (standard quintessence). In other words, the Big Crunch Singularity is delayed for $\xi=0.1$. In case of $\xi=0.5$, the slow-contraction regime does not exist any more, an accelerated expansion is obtained for longer period, and finally the universe gets de-Sitter expansion in distant future, see Fig. 3. Therefore, the Big Crunch Singularity is redundant for large values of non-minimal coupling (e.g. $\xi=0.5$) in NMC scalar field model. The numerical evolution of scale factor $a(t)$, Hubble parameter $H(t)$ and deceleration parameter $q(t)$ of NMC scalar field model are well approximated by $\Lambda$CDM as shown in Fig. 3, where we get an accelerated expansion at current epoch and de-Sitter expansion in far future.
@@ III Phase Space Analysis: Stationary points and their stability
The equations of motion (2)--(5) are reformulated in new variables as a system of differential equations of first order.
@@ III.1 Cosmological Importance
Points $A_{3}$ and $A_{4}$ (Stable Attractor Solution): The fixed points $A_{3}$ and $A_{4}$ are stable attractors, which behave as a stable node with eigenvalues given by equations (III) and (III), respectively. It further shows that the solutions exhibit damped oscillatory behavior along the real parts of the eigenvalues with a stable attractior to the fixed point. At these points, the universe shows accelerating behavior with $\Omega_{\phi}=1$, meaning the scalar field dominates the energy density, while $\Omega_{m}=0$ indicates no matter content. The EoS is $w_{eff}=w_{\phi}=-1$, corresponding to a cosmological constant-like behavior, which is consistent with an accelerated expansion of the universe. The phase portraits provide crucial insight into the attractor behavior for both the points. As shown in Fig. 4, all trajectories approach the attractor point, which validates the stability of the fixed point and its role in describing the late-time behavior of the universe, particularly in the context of dark energy dominance. The stable node is typical of models where dark energy dominates at late-time. It corresponds to a cosmological constant or quintessence field behavior, where the universe enters a phase of accelerated expansion.
@@ IV Equations for Data analysis
In this section, we reconstruct the first-order dynamical equations that facilitate the evaluation of the Hubble parameter in our model. In the previous section, we analyzed the qualitative behavior of the system using the dynamical systems framework. However, the equations expressed in that form exhibit numerical instability. Such instability is common in highly non-linear models; therefore, reformulating the system in a slightly different manner can significantly enhance numerical stability. It is important to note that the autonomous system presented earlier can still be used to derive constraints on the model parameters. In the previous works, the parameters for the $k-$essence and quintessence scalar field model have been constrained using the dynamical system analysis framework [69, 70, 71]. The reformulation presented here does not alter the physical interpretation or the parameter constraints of the system. Therefore, the redefined variables yield equivalent information. The redefined dimensionless dynamical variables are
@@ V Data Analysis
  • BAO data: The Baryonic Acoustic Oscillation (BAO) dataset, referred to as DESBAO, consists of 7 data points from the Dark Energy Spectroscopic Instrument (DESI) Release II [78], which is an updated and enhanced version of DR1 [77, 80, 81]. We derive the constraints using the full covariance matrix provided in [78]. The relevant observables include $(D_{M},D_{H},D_{V},r_{d})$, corresponding to the comoving angular distance, comoving Hubble distance, spherically averaged distance, and the sound horizon at the drag epoch, respectively. For mathematical definitions of these parameters, we refer readers to Refs. [79, 77]. In this study, we treat $r_{d}$ as a free parameter, which we constrain using various combinations of datasets. Specifically, we consider two distinct combinations:
@@
    where $k$ is the number of model parameters, $N$ is the total number of data points, and $\mathcal{L}_{\rm max}$ is the maximum likelihood. To assess model preference, we compute the difference in information criteria (IC) between the current model and $\Lambda$CDM:
@@ VI Results and Discussion
Additionally, we present the correlation matrix in Fig. 6, generated using Seaborn python package, utilizing the sample chains obtain from MCMC [90]. The matrix is constructed by combining samples from both datasets, as the parameter correlations are found to be qualitatively similar across them. In this matrix, higher positive (or negative) values indicate stronger correlations between the parameters. A positive correlation implies that as one parameter increases, the other tends to increase proportionally, whereas a negative correlation indicates inverse proportionality. For example, $H_{0}$ and $r_{d}$ exhibit a strong negative correlation, implying that an increase in $H_{0}$ corresponds to a decrease in $r_{d}$, and vice versa. A similar inverse correlation is observed between $\Omega_{0m}$ and $r_{d}$. In contrast, the correlation between $\Omega_{0m}$ and $H_{0}$ is weak and also negative. Among the model parameters, $\alpha$ and $\beta$, which define the structure of the potential, are strongly anti-correlated. Furthermore, $\alpha$ is positively correlated with $z_{0}$, the initial value of the field variable $z=\kappa\phi$, while showing negligible correlation with $\beta$. The parameter $A_{0}$, on the other hand, shows negative correlation with both $\alpha$ and $\beta$, suggesting a suppressive effect on the potential parameters.
@@
The qualitative behavior of the models is illustrated in Fig. 7 as a function of redshift $z$, where we plot the fractional energy densities of the scalar field and matter fluid, the EoS of the field, and the effective EoS of the system, using the best-fit parameter values obtained from the MCMC analysis. The evolution of the field energy density indicates that it begins to dominate near the current epoch, with the corresponding EoS of the field approaching $\omega_{\phi}\approx-1$, while the effective EoS satisfies $\omega_{eff}<-0.5$, mimicking the behavior of the $\Lambda$CDM model. At present, the matter density is approximately $\Omega_{m}\sim 0.3$, and it increases at higher redshifts, reflecting the expected matter-dominated era where the effective EoS tends toward zero. As the system is extrapolated into the asymptotic past, the effective EoS rises toward $\omega_{eff}\to 1$, signaling a stiff matter-like behavior, with $\Omega_{eff}$ correspondingly dominating. It is important to note that, since radiation has not been included in this system, the onset of stiff matter behavior occurs at intermediate redshifts for specific parameter configurations. Inclusion of a radiation component would shift this stiff phase further back into the early universe. However, given that the observational data used in this study are largely limited to redshifts $z\lesssim 3$, the radiation density does not significantly affect our late-time cosmological analysis.
@@
We assess the statistical performance of the proposed models using information criteria and compare them with the standard $\Lambda$CDM model as shown in Fig. 9. In addition, we compute the reduced chi-square statistic, defined as $\chi_{\rm rd}^{2}\equiv\chi_{\rm min}^{2}/\nu$, where $\nu$ represents the number of observational data points minus the number of fitted parameters. This metric provides an important measure of goodness-of-fit. A value of $\chi_{\rm rd}^{2}\ll 1$ typically indicates model overfitting, whereas $\chi_{\rm rd}^{2}\gg 1$ suggests a poor fit. A value near unity implies that the model fits the data well. Our results show that the proposed models yield $\chi_{\rm rd}^{2}\sim 1$, consistent with that of the $\Lambda$CDM model, indicating excellent agreement with the observational data. Furthermore, we find that the current models yield a $\Delta\rm AIC$ of approximately 5, suggesting moderate support for the models relative to $\Lambda$CDM, though not sufficient for a strong preference. Conversely, the $\Delta\rm BIC>10$ indicates weak support for the current models when penalizing model complexity more strongly. Among the two considered cases, the model with $\xi=0.1$ exhibits a slightly better performance in both AIC and BIC metrics compared to the model with $\xi=0.05$, thus indicating a mild statistical preference.